# Towards a Theoretical Understanding of Memorization in Diffusion Models

### v1

@@ 1 Introduction
The diffusion probabilistic models (DPMs) (Ho et al., 2020; Sohl-Dickstein et al., 2015; Song & Ermon, 2019) is a family of powerful generative models that learn the distribution of a dataset by first gradually destroying the structure of the data through an iterative forward diffusion process and then restoring the data structure via a reverse diffusion process. Due to their outstanding capability in capturing data distribution, DPMs have become the foundation models for many pioneering Generative Artificial Intelligence (GenAI) products such as Stable Diffusion (Rombach et al., 2022), DALL-E 3 (Betker et al., ), and Sora (Brooks et al., 2024). Despite the widespread adoption of DPMs, a potential risk they face is data memorization, i.e., the risk of memorizing a certain proportion of the raw training samples. This could result in the generation of memorized (rather than new) samples via direct copying, causing data leakage, privacy breaches, or copyright infringement, as highlighted in the literature (Somepalli et al., 2022; 2023; Asay, 2020; Cooper & Grimmelmann, 2024). A current case argues that Stable Diffusion is a 21st-century collage tool, remixing the copyrighted creations of countless artists whose works were included in the training data(Butterick, 2023). Furthermore, data memorization also gives rise to data extraction attacks, which is one type of privacy attacks that attempt to extract the training data from a well-trained model. Notably, recent work by Carlini et al. (2023) demonstrated the feasibility of extracting training data samples from DPMs like Stable Diffusion (Rombach et al., 2022), revealing the potential dangers associated with these models.
Several works have investigated the data memorization phenomenon in diffusion models. For example, it has been observed that there exists a strong correlation between training data memorization and conditional DPMs, i.e., being conditional is more prone to memorization (Somepalli et al., 2023). Gu et al. (2023) investigates the influential factors on memorization behaviors via a comprehensive set of experiments. They found that conditioning on random-labeled data can significantly trigger memorization, and unconditional models memorize much less training data. Though inspiring, these understandings are primarily empirical. Moreover, without a unified theoretical understanding of memorization in both conditional and unconditional DPMs, it is extremely difficult to design an effective data extraction method for unconditional DPMs.
@@
  • We evaluate the effectiveness of SIDE on CIFAR-10 and various scales of CelebA datasets, and show that, on average, it can outperform the baseline method proposed by Carlini et al. (2023) by more than 50%.
@@ 2 Related Work
Diffusion Probabilistic Models  DPMs (Sohl-Dickstein et al., 2015) such as Stable Diffusion (Rombach et al., 2022), DALL-E 3 (Betker et al., ), Sora (Brooks et al., 2024), Runway (Rombach et al., 2022), and Imagen (Saharia et al., 2022) have achieved state-of-the-art performance in image/video generation across a wide range of benchmarks (Dhariwal & Nichol, 2021). These models can be viewed from two perspectives. The first is score matching (Song & Ermon, 2019), where diffusion models learn the gradient of the image distribution (Song et al., 2020). The second perspective involves denoising DPMs (Ho et al., 2020), which add Gaussian noise at various time steps to clean images and train models to denoise them. To conditionally sample from diffusion models, (Dhariwal & Nichol, 2021) utilizes a classifier to guide the denoising process at each sampling step. Additionally, (Ho & Salimans, 2022) introduces classifier-free guidance for conditional data sampling using DPMs.
@@
Memorization in Diffusion Models  Early exploration of memorization in large models was focused on language models (Carlini et al., 2022; Jagielski et al., 2022), which has motivated more in-depth research on image-generation DPMs (Somepalli et al., 2023; Gu et al., 2023). Notably, a recent research on image-generation DPMs. Somepalli et al. (2022) found that 0.5-2% of generated images duplicate training samples, a result concurrently reported by Carlini et al. (2023) in broader experiments for both conditional and unconditional diffusion models. Further studies Somepalli et al. (2023) and Gu et al. (2023) linked memorization to model conditioning, showing that conditional models are more prone to memorization. To address this issue, anti-memorization guidance has been proposed to mitigate memorization during the sampling process Chen et al. (2024a). So far, the motivated data extraction attack or defend methods from these empirical understandings are mostly for conditional DPMs Carlini et al. (2023); Webster (2023), and studies have shown that extracting training data from unconditional DPMs can be much more challenging than that on conditional DPMs Gu et al. (2023). Although the empirical understandings are inspiring, a theoretical understanding of the memorization behaviors of DPMs is missing from the current literature. A recent attempt by Ross et al. (2024) uses the local intrinsic dimension (LID) metric to characterize memorization in DPMs. However, the finding that lower LID leads to more memorization is only verified on a toy 1-dimensional dataset and only a few generated images, and it fails to address the data extraction challenge on unconditional DPMs. This challenge is exemplified by the brute-force methods currently employed, as discussed in (Somepalli et al., 2022; Carlini et al., 2023), highlighting the inadequacies of existing approaches for extracting data from unconditional diffusion models. In response to these challenges, we introduce a theoretical framework to characterize memorization in DPMs, which further motivates a novel data extraction method for unconditional DPMs.
@@ 3.1 Memorization Metric
A straightforward choice for $q({\bm{x}},{\bm{x}}_{i},\epsilon)$ is the Dirac delta function centered at training data point ${\bm{x}}_{i}$: $q({\bm{x}},{\bm{x}}_{i},\epsilon)=\delta({\bm{x}}-{\bm{x}}_{i})$. However, this would make Eq. (1) uncomputable as the Dirac delta function is zero beyond the $\epsilon$-neighborhood. Alternatively, we could use the Gaussian distribution with a covariance matrix $\epsilon\textit{I}$ (I is the identity matrix) for $q({\bm{x}},{\bm{x}}_{i},\epsilon)$:
@@
Note that in Eq. (1), a smaller value of ${\mathcal{M}}_{point}({\mathcal{D}};\theta)$ indicates more overlap between the two distributions and thus more memorization. As $\epsilon$ tends to 00, the measured memorization becomes more accurate, with the limit characterizing the intrinsic memorization capability of model $p_{\theta}$.
@@
We first build the relationship map ${\mathcal{M}}_{point}$ into latent space, and then break down Eq. (4) into separate components and derive each component separately. So, the Eq. (4) can be simplified to $\frac{1}{2\epsilon}$. The detailed proof can be found in Appendix A.2. Note that the memorization effect studied in this work refers specifically to the pointwise memorization ${\mathcal{M}}_{point}$.
@@ 3.2 Theoretical Framework
The concept of informative labels has been previously discussed as class labels (Gu et al., 2023). In this work, we introduce a more general definition that encompasses both class labels and random labels as special cases. We define an informative label as follows.
@@
Here, the labels are not limited to the conventional class labels; they can also be text captions, features, or cluster information that can be used to group the training samples into subsets. The above definition states that an information label should be able to differentiate a subset of samples from others. An extreme case is that all samples have the same label; in this case, the label is not informative. According to our definition, class-wise and random labels are special cases of informative labels. Informative labels can be either explicit like class/random labels and text captions, or implicit like silent features or clusters. Next, we will explore the correlation between informative labels and the clustering effect in the latent space of a generative model with an encoder and decoder. In diffusion models, the encoder represents the forward diffusion process, while the decoder represents the reverse process. Notably, the latent space can be defined at earlier timesteps according to (Ho et al., 2020).
@@
Given suffix training on ${\mathcal{D}}$, the generative model $p_{\theta}$ converges to an optimal solution for Eq. (6) : $\theta^{*}=\arg\min_{\theta}-\sum_{{\bm{x}}_{\mathrm{i}}\in\mathcal{D}}\log p_{\theta}({\bm{x}}_{i}|y_{i})$.
@@
Proposition 1 describes two properties of the learned latent space driven by informative labels: 1) tighter clustering as defined in Eq. (7) and 2) reduced variance as defined in Eq. (8). Tighter clustering means that the data samples associated with the same informative label are more closely clustered, which allows the model to more effectively capture and memorize the specific features and patterns relevant to those labels. This proximity in the latent space enhances the model’s ability to recall memorized samples during generation, as the representations are organized around distinct means. Additionally, reduced variance in these clustered representations leads to greater stability, ensuring that the model can consistently reproduce memorized outputs from the clustered latent codes. Based on this understanding, we formalize the relationship between informative labels and memorization in conditional DPMs via the following theorem.
@@
The proof can be found in Appendix A.4. Theorem 2 states that any form of information labels can incur more memorization in DPMs, including conventional class labels and random labels. As shown in our proof, the informative nature of a label reduces the entropy (or uncertainty) of the data distribution conditioned on that label, leading to a more focused and memorable data representation. This explains the two empirical observations made in existing works Gu et al. (2023): 1) conditional DPMs are more prone to memorization and 2) even random labels can lead to more memorization. It also explains the findings that unconditional models do not replicate data and that text conditioning increases memorization Somepalli et al. (2022); Chen et al. (2024a).
@@ 4.1 Constructing Implicit Informative Labels
We can use the classifier that generates the implicit labels to approximate the gradient in Eq. (10). However, a known challenge associated with neural network classifiers is their tendency towards miscalibration (Guo et al., 2017). Specifically, the classifier could be overconfident or underconfident about its predictions. To mitigate the potential impact of miscalibration on the sampling process, we introduce a hyperparameter $\lambda$ to calibrate the classifier’s probability output on the diffusion path using power prior as follows:
@@ 4.2 Time-dependent Classifier
In Eq. (151), the classifier is denoted by $p_{\theta}^{t}(y|{\bm{x}})$ with $t$ being the timestep, implying its time-dependent nature. However, we do not have a time-dependent classifier but only a time-independent classifier. To address this problem, we propose a method named Time-Dependent Knowledge Distillation (TDKD) to train a time-dependent classifier. The distillation process is illustrated in Figure 2. TDKD equips the classifier with time-dependent guidance during sampling. It operates in two steps: first, the network architecture is adjusted to accommodate time-dependent inputs, and the structure of the time-dependent module and modification are illustrated in Appendix C; second, a synthetic dataset and pseudo labels are created to facilitate knowledge distillation from the normal classifier to its time-dependent counterpart.
As the original training dataset is unknown, we employ the target DPM to generate a synthetic dataset, following the generative data augmentation techniques Chen et al. (2023; 2024b). We then use the normal classifier trained on the original dataset to generate pseudo labels for the synthetic dataset. Finally, we train a time-dependent classifier on the labelled synthetic dataset. The objective of this training is defined as following:
@@ 5.1 Experimental Setup
We evaluate the effectiveness of our method on three datasets: 1) CelebA-HQ-Face-Identity (CelebA-HQ-FI) (Na et al., 2022) which consists of 5478 images, 2) a subset of the CelebA (CelebA-25000) (Liu et al., 2015) which contains 25,000 images and 2) CIFAR-10 containing 50,000 images. All the images are normalized to [-1,1]. We use the AdamW optimizer (Loshchilov & Hutter, 2019) with a learning rate of 1e-4 to train the time-dependent classifier. We train denoising DPMs with a discrete denoising scheduler (DDIM (Song et al., 2021)) using the HuggingFace implementation (von Platen et al., 2022). All DPMs are trained with batch size 64. We train the models for 258k steps ($\approx$ 3000 epochs) on CelebA-HQ-FI, 390k steps ($\approx$ 1000 epochs) on CelebA-25000 and 1600k steps ($\approx$ 2048 epochs) on CIFAR-10, respectively. We use ResNet34 (He et al., 2015) as the normal classifier.
@@ 5.2 Performance Metrics
Determining whether a generated image is a memorized copy of a training image is difficult, as $L_{p}$ distances in the pixel space are ineffective. Previous research addresses this by using the 95th percentile Self-Supervised Descriptor for Image Copy Detection (SSCD) score for image copy detection (Somepalli et al., 2022; Gu et al., 2023). However, the 95th percentile SSCD score has three limitations: 1) it does not measure the uniqueness of memorized images; 2) it may underestimate the number of memorized samples when cut at the 95th percentile; and 3) it does not account for different types of memorization. Here, we propose two new memorization scores to solve these issues: 1) Average Memorization Score (AMS) and 2) Unique Memorization Score (UMS metrics).
@@
where ${{\mathcal{D}}}_{\text{gen}}$ is the generated dataset, ${\mathcal{D}}_{\text{train}}$ is the training dataset, and $\alpha$, $\beta$ are thresholds for image similarity scoring. $\mathcal{F}({\bm{x}}_{i},{\mathcal{D}}_{\text{train}},\alpha,\beta)$ serves as a binary check for whether any training sample meets the similarity/distance criteria. $\phi({\bm{x}}_{i},{\mathcal{D}}_{\text{train}},\alpha,\beta)$ provides the specific indices of those training samples that meet the similarity/distance criteria. Mathematically, we can represent $\mathcal{F}(x_{i},{\mathcal{D}}_{\text{train}},\alpha,\beta)=\mathds{1}[\max_{x_{j}\in{\mathcal{D}}_{\text{train}}}\gamma(x_{i},x_{j})\geq\alpha\;\&\;\gamma(x_{i},x_{j})\leq\beta]$ $\varphi(x_{i},{\mathcal{D}}_{\text{train}},\alpha,\beta)=\{j:x_{j}\in{\mathcal{D}}_{\text{train}},\gamma(x_{i},x_{j})\geq\alpha\;\&\;\gamma(x_{i},x_{j})\leq\beta\}$. $\gamma$ represents the similarity/distance function. For low-resolution datasets, we use the normalized $L_{2}$ distance as $\gamma$ following Carlini et al. (2023), while for high-resolution datasets, we use the SSCD score as $\gamma$. In our experiments, the thresholds for SSCD are set to $\alpha=0.4$ and $\beta=0.5$ for low similarity, $\alpha=0.5$ and $\beta=0.6$ for mid similarity, and $\alpha=0.6$ and $\beta=1.0$ for high similarity. The thresholds for the normalized $L_{2}$ ditance (Carlini et al., 2023) are set to $\alpha=1.5$ and $\beta=10$ for low similarity, $\alpha=1.4$ and $\beta=1.5$ for mid similarity, and $\alpha=1.35$ and $\beta=1.4$ for high similarity.
@@
While Carlini et al. (2023); Chen et al. (2024a) introduced metrics similar to AMS and UMS, they did not account for varying levels of similarity, which is essential for assessing different types of copyright infringement, such as character or style copying (Lee et al., 2023; Sag, 2023; Sobel, 2023). Additionally, the UMS considers the number of generated images $N_{G}$, which was overlooked in (Carlini et al., 2023). The significance of $N_{G}$ lies in its non-linear impact on the UMS (see Appendix B for proof), indicating that UMS scores should not be compared across different values of $N_{G}$.

### v2

# iclr2025_conference.bib

@@ 1 Introduction
The diffusion probabilistic models (DPMs) ho2020denoising; sohl2015deep; song2019generative is a family of powerful generative models that learn the distribution of a dataset by first gradually destroying the structure of the data through an iterative forward diffusion process and then restoring the data structure via a reverse diffusion process. Due to their outstanding capability in capturing data distribution, DPMs have become the foundation models for many pioneering Generative Artificial Intelligence (GenAI) products such as Stable Diffusion rombach2022high, DALL-E 3 BetkerImprovingIG, and Sora videoworldsimulators2024. Despite the widespread adoption of DPMs, a potential risk they face is data memorization, i.e., the risk of memorizing a certain proportion of the raw training samples. This could result in the generation of memorized (rather than new) samples via direct copying, causing data leakage, privacy breaches, or copyright infringement, as highlighted in the literature Somepalli2022DiffusionAO; somepalli2023understanding; asay2020independent; cooper2024files. A current case argues that Stable Diffusion is a 21st-century collage tool, remixing the copyrighted creations of countless artists whose works were included in the training databutterick2023stable. Furthermore, data memorization also gives rise to data extraction attacks, which is one type of privacy attacks that attempt to extract the training data from a well-trained model. Notably, recent work by carlini2023extracting demonstrated the feasibility of extracting training data samples from DPMs like Stable Diffusion rombach2022high, revealing the potential dangers associated with these models.
Several works have investigated the data memorization phenomenon in diffusion models. For example, it has been observed that there exists a strong correlation between training data memorization and conditional DPMs, i.e., being conditional is more prone to memorization somepalli2023understanding. Gu2023OnMI investigates the influential factors on memorization behaviors via a comprehensive set of experiments. They found that conditioning on random-labeled data can significantly trigger memorization, and unconditional models memorize much less training data. Though inspiring, these understandings are primarily empirical. Moreover, without a unified theoretical understanding of memorization in both conditional and unconditional DPMs, it is extremely difficult to design an effective data extraction method for unconditional DPMs.
@@
  • We evaluate the effectiveness of SIDE on CIFAR-10 and various scales of CelebA datasets, and show that, on average, it can outperform the baseline method proposed by carlini2023extracting by more than 50%.
@@ 2 Related Work
Diffusion Probabilistic Models  DPMs sohl2015deep such as Stable Diffusion rombach2022high, DALL-E 3 BetkerImprovingIG, Sora videoworldsimulators2024, Runway rombach2022high, and Imagen saharia2022photorealistic have achieved state-of-the-art performance in image/video generation across a wide range of benchmarks dhariwal2021diffusion. These models can be viewed from two perspectives. The first is score matching song2019generative, where diffusion models learn the gradient of the image distribution song2020score. The second perspective involves denoising DPMs ho2020denoising, which add Gaussian noise at various time steps to clean images and train models to denoise them. To conditionally sample from diffusion models, dhariwal2021diffusion utilizes a classifier to guide the denoising process at each sampling step. Additionally, ho2022classifier introduces classifier-free guidance for conditional data sampling using DPMs.
@@
Memorization in Diffusion Models  Early exploration of memorization in large models was focused on language models carlini2022quantifying; jagielski2022measuring, which has motivated more in-depth research on image-generation DPMs somepalli2023understanding; Gu2023OnMI. Notably, a recent research on image-generation DPMs. Somepalli2022DiffusionAO found that 0.5-2% of generated images duplicate training samples, a result concurrently reported by carlini2023extracting in broader experiments for both conditional and unconditional diffusion models. Further studies somepalli2023understanding and Gu2023OnMI linked memorization to model conditioning, showing that conditional models are more prone to memorization. To address this issue, anti-memorization guidance has been proposed to mitigate memorization during the sampling process chen2024towards. So far, the motivated data extraction attack or defend methods from these empirical understandings are mostly for conditional DPMs carlini2023extracting; webster2023reproducible, and studies have shown that extracting training data from unconditional DPMs can be much more challenging than that on conditional DPMs Gu2023OnMI. Although the empirical understandings are inspiring, a theoretical understanding of the memorization behaviors of DPMs is missing from the current literature. A recent attempt by ross2024a uses the local intrinsic dimension (LID) metric to characterize memorization in DPMs. However, the finding that lower LID leads to more memorization is only verified on a toy 1-dimensional dataset and only a few generated images, and it fails to address the data extraction challenge on unconditional DPMs. This challenge is exemplified by the brute-force methods currently employed, as discussed in Somepalli2022DiffusionAO; carlini2023extracting, highlighting the inadequacies of existing approaches for extracting data from unconditional diffusion models. In response to these challenges, we introduce a theoretical framework to characterize memorization in DPMs, which further motivates a novel data extraction method for unconditional DPMs.
@@ 3.1 Memorization Metric
A straightforward choice for $q({\bm{x}},{\bm{x}}_{i},\epsilon)$ is the Dirac delta function centered at training data point ${\bm{x}}_{i}$: $q({\bm{x}},{\bm{x}}_{i},\epsilon)=\delta({\bm{x}}-{\bm{x}}_{i})$. However, this would make equation (1) uncomputable as the Dirac delta function is zero beyond the $\epsilon$-neighborhood. Alternatively, we could use the Gaussian distribution with a covariance matrix $\epsilon\textit{I}$ (I is the identity matrix) for $q({\bm{x}},{\bm{x}}_{i},\epsilon)$:
@@
Note that in equation (1), a smaller value of ${\mathcal{M}}_{point}({\mathcal{D}};\theta)$ indicates more overlap between the two distributions and thus more memorization. As $\epsilon$ tends to 00, the measured memorization becomes more accurate, with the limit characterizing the intrinsic memorization capability of model $p_{\theta}$.
@@
We first build the relationship map ${\mathcal{M}}_{point}$ into latent space, and then break down equation (4) into separate components and derive each component separately. So, the equation (4) can be simplified to $\frac{1}{2\epsilon}$. The detailed proof can be found in Appendix A.2. Note that the memorization effect studied in this work refers specifically to the pointwise memorization ${\mathcal{M}}_{point}$.
@@ 3.2 Theoretical Framework
The concept of informative labels has been previously discussed as class labels Gu2023OnMI. In this work, we introduce a more general definition that encompasses both class labels and random labels as special cases. We define an informative label as follows.
@@
Here, the labels are not limited to the conventional class labels; they can also be text captions, features, or cluster information that can be used to group the training samples into subsets. The above definition states that an information label should be able to differentiate a subset of samples from others. An extreme case is that all samples have the same label; in this case, the label is not informative. According to our definition, class-wise and random labels are special cases of informative labels. Informative labels can be either explicit like class/random labels and text captions, or implicit like silent features or clusters. Next, we will explore the correlation between informative labels and the clustering effect in the latent space of a generative model with an encoder and decoder. In diffusion models, the encoder represents the forward diffusion process, while the decoder represents the reverse process. Notably, the latent space can be defined at earlier timesteps according to ho2020denoising.
@@
Given suffix training on ${\mathcal{D}}$, the generative model $p_{\theta}$ converges to an optimal solution for equation (6) : $\theta^{*}=\arg\min_{\theta}-\sum_{{\bm{x}}_{\mathrm{i}}\in\mathcal{D}}\log p_{\theta}({\bm{x}}_{i}|y_{i})$.
@@
Proposition 1 describes two properties of the learned latent space driven by informative labels: 1) tighter clustering as defined in equation (7) and 2) reduced variance as defined in equation (8). Tighter clustering means that the data samples associated with the same informative label are more closely clustered, which allows the model to more effectively capture and memorize the specific features and patterns relevant to those labels. This proximity in the latent space enhances the model’s ability to recall memorized samples during generation, as the representations are organized around distinct means. Additionally, reduced variance in these clustered representations leads to greater stability, ensuring that the model can consistently reproduce memorized outputs from the clustered latent codes. Based on this understanding, we formalize the relationship between informative labels and memorization in conditional DPMs via the following theorem.
@@
The proof can be found in Appendix A.4. Theorem 2 states that any form of information labels can incur more memorization in DPMs, including conventional class labels and random labels. As shown in our proof, the informative nature of a label reduces the entropy (or uncertainty) of the data distribution conditioned on that label, leading to a more focused and memorable data representation. This explains the two empirical observations made in existing works Gu2023OnMI: 1) conditional DPMs are more prone to memorization and 2) even random labels can lead to more memorization. It also explains the findings that unconditional models do not replicate data and that text conditioning increases memorization Somepalli2022DiffusionAO; chen2024towards.
@@ 4.1 Constructing Implicit Informative Labels
We can use the classifier that generates the implicit labels to approximate the gradient in equation (10). However, a known challenge associated with neural network classifiers is their tendency towards miscalibration guo2017calibration. Specifically, the classifier could be overconfident or underconfident about its predictions. To mitigate the potential impact of miscalibration on the sampling process, we introduce a hyperparameter $\lambda$ to calibrate the classifier’s probability output on the diffusion path using power prior as follows:
@@ 4.2 Time-dependent Classifier
In equation (151), the classifier is denoted by $p_{\theta}^{t}(y|{\bm{x}})$ with $t$ being the timestep, implying its time-dependent nature. However, we do not have a time-dependent classifier but only a time-independent classifier. To address this problem, we propose a method named Time-Dependent Knowledge Distillation (TDKD) to train a time-dependent classifier. The distillation process is illustrated in Figure 2. TDKD equips the classifier with time-dependent guidance during sampling. It operates in two steps: first, the network architecture is adjusted to accommodate time-dependent inputs, and the structure of the time-dependent module and modification are illustrated in Appendix C; second, a synthetic dataset and pseudo labels are created to facilitate knowledge distillation from the normal classifier to its time-dependent counterpart.
As the original training dataset is unknown, we employ the target DPM to generate a synthetic dataset, following the generative data augmentation techniques chen2023data; chen2024comprehensive. We then use the normal classifier trained on the original dataset to generate pseudo labels for the synthetic dataset. Finally, we train a time-dependent classifier on the labelled synthetic dataset. The objective of this training is defined as following:
@@ 5.1 Experimental Setup
We evaluate the effectiveness of our method on three datasets: 1) CelebA-HQ-Face-Identity (CelebA-HQ-FI) na2022unrestricted which consists of 5478 images, 2) a subset of the CelebA (CelebA-25000) liu2015faceattributes which contains 25,000 images and 2) CIFAR-10 containing 50,000 images. All the images are normalized to [-1,1]. We use the AdamW optimizer loshchilov2018decoupled with a learning rate of 1e-4 to train the time-dependent classifier. We train denoising DPMs with a discrete denoising scheduler (DDIM song2021denoising) using the HuggingFace implementation diffusers. All DPMs are trained with batch size 64. We train the models for 258k steps ($\approx$ 3000 epochs) on CelebA-HQ-FI, 390k steps ($\approx$ 1000 epochs) on CelebA-25000 and 1600k steps ($\approx$ 2048 epochs) on CIFAR-10, respectively. We use ResNet34 He2015DeepRL as the normal classifier.
@@ 5.2 Performance Metrics
Determining whether a generated image is a memorized copy of a training image is difficult, as $L_{p}$ distances in the pixel space are ineffective. Previous research addresses this by using the 95th percentile Self-Supervised Descriptor for Image Copy Detection (SSCD) score for image copy detection Somepalli2022DiffusionAO; Gu2023OnMI. However, the 95th percentile SSCD score has three limitations: 1) it does not measure the uniqueness of memorized images; 2) it may underestimate the number of memorized samples when cut at the 95th percentile; and 3) it does not account for different types of memorization. Here, we propose two new memorization scores to solve these issues: 1) Average Memorization Score (AMS) and 2) Unique Memorization Score (UMS metrics).
@@
where ${{\mathcal{D}}}_{\text{gen}}$ is the generated dataset, ${\mathcal{D}}_{\text{train}}$ is the training dataset, and $\alpha$, $\beta$ are thresholds for image similarity scoring. $\mathcal{F}({\bm{x}}_{i},{\mathcal{D}}_{\text{train}},\alpha,\beta)$ serves as a binary check for whether any training sample meets the similarity/distance criteria. $\phi({\bm{x}}_{i},{\mathcal{D}}_{\text{train}},\alpha,\beta)$ provides the specific indices of those training samples that meet the similarity/distance criteria. Mathematically, we can represent $\mathcal{F}(x_{i},{\mathcal{D}}_{\text{train}},\alpha,\beta)=\mathbb{1}[\max_{x_{j}\in{\mathcal{D}}_{\text{train}}}\gamma(x_{i},x_{j})\geq\alpha\;\&\;\gamma(x_{i},x_{j})\leq\beta]$ $\varphi(x_{i},{\mathcal{D}}_{\text{train}},\alpha,\beta)=\{j:x_{j}\in{\mathcal{D}}_{\text{train}},\gamma(x_{i},x_{j})\geq\alpha\;\&\;\gamma(x_{i},x_{j})\leq\beta\}$. $\gamma$ represents the similarity/distance function. For low-resolution datasets, we use the normalized $L_{2}$ distance as $\gamma$ following carlini2023extracting, while for high-resolution datasets, we use the SSCD score as $\gamma$. In our experiments, the thresholds for SSCD are set to $\alpha=0.4$ and $\beta=0.5$ for low similarity, $\alpha=0.5$ and $\beta=0.6$ for mid similarity, and $\alpha=0.6$ and $\beta=1.0$ for high similarity. The thresholds for the normalized $L_{2}$ ditance carlini2023extracting are set to $\alpha=1.5$ and $\beta=10$ for low similarity, $\alpha=1.4$ and $\beta=1.5$ for mid similarity, and $\alpha=1.35$ and $\beta=1.4$ for high similarity.
@@
While carlini2023extracting; chen2024towards introduced metrics similar to AMS and UMS, they did not account for varying levels of similarity, which is essential for assessing different types of copyright infringement, such as character or style copying lee2023talkin; sag2023copyright; sobel2023elements. Additionally, the UMS considers the number of generated images $N_{G}$, which was overlooked in carlini2023extracting. The significance of $N_{G}$ lies in its non-linear impact on the UMS (see Appendix B for proof), indicating that UMS scores should not be compared across different values of $N_{G}$.